How much potassium chlorate is neede to get enough oxygen for completely burning of carbon mono oxide at NTP?

Grade:11

1 Answers

Vikas TU
6 years ago
2KClO3 ==> 2KCl + 3O2 
You require what number of mols O2? That is 
n = PV/RT. Substitute and tackle for n = mols O2.' 
At ntp , p = 1 atm , t = 293 k , 
R = 259.8 for oxygen 
2 = 1 x v/293 x 259.8 = 
2 = v/76,121.4 
V= 76,121.4 x 2 
152.8 l 
1 Mole of oxygen has 24 volume 
X mole has 152.8 , x = 24/152.8 = 0.157 moles of oxygen .